A company has two products: A and B. It uses activity-based costing and has prepared the following analysis showing budgetedMultiple Choice $3.49 $2.25 $10.28 $12.44 $1.77
Answer #1

Answer is $2.25

Calculated as

Calculation of Activity overhead rate
Activity Activity
Activity Cost Pool Cost Product A Product B Total Activity = A+B Overhead Rate = Cost/Cost Driver
1 97000 4000 3800 7800 $                                                        12.44
2 72000 5500 6500 12000 $                                                          6.00
3 113000 3500 6250 9750 $                                                        11.59
Calculated of Overhead Cost of Product B
Activity Cost Pool Overhead Rate Activity
1 $    12.44 x 3800 = $                                               47,256.41
2 $       6.00 x 6500 = $                                               39,000.00
3 $    11.59 x 6250 = $                                               72,435.90
Total overhead cost of 70550 units $                                             158,692.31
Per unit overhead cost of Product B = $158692.31/70550= $       2.25

Hit Thumbs up if satisfied

Have any query mention in comment section please

Thank you